OpenStudy (anonymous):

375ml=π r² × 12cm How to get r? Please explain step by step

OpenStudy (bradely):

375ml=π r² × 12cm 375 cm^3 =π r² × 12cm r² =375/12π r² =9.95 r=3.15 cm post on: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

How to get 3.15cm from r² =9.95?

OpenStudy (anonymous):

square root the r

OpenStudy (anonymous):

Thank you so much for your help!